OpenStudy (anonymous):

Solve the equation by graphing: 3x2 + 9x = 12. A. {-4, 1} B. {-6, 2} C. {-3, 4} D. No solution Solve the equation by graphing: 3x2 + 9x = 12. A. {-4, 1} B. {-6, 2} C. {-3, 4} D. No solution @Mathematics

OpenStudy (anonymous):

What are you looking for? Vertex? Point on the parabola? Intercept?

hero (hero):

3x^2+9x-12 = 0 3x^2-3x+12x-12=0 3x(x-1)-12(x-1)=0 (x-1)(3x-12)=0 x=-4,1

OpenStudy (anonymous):

Oh those aren't coordinates, they're solution sets. Well still, I can't really graph it for you, but I can factor it. 3(x^2+3x-4) 3(x+4)(x-1) x=-4,1